# A Room-Temperature Solid-State Maser Amplifier

### v1

@@ Abstract
Masers once represented the state-of-the-art in low noise microwave amplification technology, but eventually became obsolete due to their need for cryogenic cooling. Masers based on solid-state spin systems perform most effectively as amplifiers, since they provide a large density of spins and can therefore operate at relatively high powers. Whilst solid-state masers oscillators have been demonstrated at room temperature, continuous-wave amplification in these systems has only ever been realized at cryogenic temperatures. Here we report on a continuous-wave solid-state maser amplifier operating at room temperature. We achieve this feat using a practical setup that includes an ensemble of nitrogen-vacancy center spins in a diamond crystal, a strong permanent magnet and simple laser diode. We describe important amplifier characteristics including gain, bandwidth, compression power and noise temperature and discuss the prospects of realizing a room-temperature near-quantum-noise-limited amplifier with this system. Finally, we show that in a different mode of operation the spins can be used to cool the system noise in an external circuit to cryogenic levels, all without the requirement for physical cooling.
@@ I Introduction
Maser amplification of short ($\sim 30~{}\mu$s) pulsed microwave signals was recently observed at room temperature using a spin system in an orgnic host [13], whilst a continuously operating diamond maser amplifier was realized at liquid nitrogen temperatures ($\sim 77$ K) [12]. However, the demonstration of a widely usable continuous-wave maser amplifier operating under ambient conditions remains a highly desirable objective.
@@
In this work we describe a continuous-wave maser amplifier based on a solid-state system that operates at room temperature. We use an ensemble of nitrogen-vacancy (NV) center spins in a bulk diamond crystal as the gain medium and couple this to a high quality factor microwave dielectric resonator. We achieve gains as high as 30 dB and a gain-bandwidth-product of up to 4.5 MHz. Our analysis shows that the intrinsic noise temperature of the spin gain medium is close to the quantum limit, whilst the dominant source of noise comes from resonator losses. We further demonstrate that the same device can be used to remove microwave noise in a circuit [16, 17, 18, 19] by acting like a matched load with an effective microwave temperature of 66 K, without the need for any cryogenic cooling. We develop a modern quantum optics theoretical model [20] to describe the maser amplifier and microwave cooler system and obtain quantitative agreement with our measurement results. Finally, we discuss relatively straight-forward improvements that can be made to our system to push its noise performance towards the quantum limit. These results demonstrate that NV spin ensembles in diamond coupled to microwave resonators constitute a system with exceptional promise for performing ultra-low noise detection of microwave signals at room temperature.
@@ II.1 Maser amplifier design
In this work we utilize a high concentration NV- diamond bulk crystal as the gain medium. The energy level structure of the NV- triplet ground state is shown in Fig. 1a. When a static magnetic field $B_{\rm 0}$ is applied parallel to one of the four equivalent NV $\langle 111\rangle$ directions (see Fig. 1b), the $m_{\rm s}=\pm 1$ states of the NV- spins Zeeman split with a gradient of ±$\pm\gamma_{\rm e}$ (with $\gamma_{\rm e}/2\pi=28$ GHz/T), offset by a zero field splitting of $D/2\pi=2.87$ GHz [25, 26]. The NV- spins can be readily polarized into the ground triplet $m_{\rm s}=0$ state by optical pumping using green ($\sim$ 520 nm) light. At a sufficiently large $B_{\rm 0}$, the $m_{\rm s}=-1$ spin state energy falls below that of the the $m_{\rm s}=0$ state and the optical pumping results in a population inversion [9] (see Fig. 1c, lower), fulfilling a key condition for implementing a maser amplifier. For smaller $B_{\rm 0}$ fields where the $m_{\rm s}=0$ state is lowest in energy (see Fig. 1c, upper), the NV- spins absorb microwaves and can thus be used to reduce the noise present in a circuit [16, 17, 18, 19].
@@ II.3.1 Gain
We proceed to operate the system as a maser amplifier by pumping the NV- spins with the laser diode at a power of $P_{\rm L}=1.61$ W, whilst simultaneously probing the system using a vector network analyzer (VNA), allowing us to measure the reflection coefficient $r^{+}$ (Eq. 2). We adjust the position of the permanent magnet poles to provide a field close to $B_{\rm 0}\approx 450$ mT, which brings the $m_{\rm s}=-1$ to $m_{s}=0$ spin transition near resonance with the dielectric resonator. We perform a fine sweep of the field using the Helmholtz coil and plot the power gain $G=20\log{|r^{+}|}$ as a function of $B_{\rm 0}$ in Fig. 3a. Away from resonance with the spins, we observe observe a dip in $G$ at $\omega_{\rm r}$, which is due to the internal resonator losses. As the spins come into resonance with the resonator, they add energy to the system via stimulated emission, which compensates the internal losses in the resonator and transforms the dip in $G$ to a peak, signalling gain. We find a single gain peak, rather than the three that would be expected due to the hyperfine coupling of the NV- electronic spin with the nitrogen nuclear spin ($I=1$) [26]. This is due to the inhomogeneity of the $B_{\rm 0}$ field supplied by the permanent magnet at the given pole separation, which broadens the spin transition and produces a Gaussian-like profile (see Supplementary Information, Section LABEL:sup:subsec:spectra).
@@ II.4 Microwave cooling
Next we adjust the permanent magnet pole positions to set $B_{\rm 0}\approx 250$ mT, bringing the $m_{\rm s}=0$ to $m_{\rm s}=+1$ spin transition in resonance with the dielectric resonator. The NV- spins, which are now polarized in the ground state of this two-level subsystem, absorb microwave photons from the resonator (including thermal photons), which are then removed from the system via optical pumping. The removal of thermal noise photons lowers the effective microwave temperature of the mode [16, 17, 18, 19], which can be transferred to an external microwave circuit.
@@
To find the cooling spin transition we perform a fine sweep of $B_{\rm 0}$ whilst probing the resonator in reflection, with the result shown in Fig. 5a. We observe avoided crossings in the measured spectrum, which is indicative of strong coupling between the resonator and hyperfine transitions of the NV- spins [18]. We note that at this magnet pole separation we are able to partially resolve the hyperfine structure of the NV- system, indicating a higher field homogeneity than observed at the maser transition. Taking a line cut (dashed line in Fig. 5a) through the middle of this feature, we observe four dips in the magnitude response (see Fig. 5b, dark blue line), which are reproduced well with our model (see Eq. 2). A fit to the magnitude ($|r^{-}|$) and phase ($\angle r^{-}$) reflection data allows us to extract the ensemble coupling strength $g_{\rm ens}$, which is plot as an inset in Fig. 5b as a function of $P_{\rm L}$. We find excellent agreement with the $g_{\rm ens}$ extracted via the gain measurements in Section II.3.1.

### v2

@@ Abstract
Masers once represented the state-of-the-art in low noise microwave amplification technology, but eventually became obsolete due to their need for cryogenic cooling. Masers based on solid-state spin systems perform most effectively as amplifiers, since they provide a large density of spins and can therefore operate at relatively high powers. Whilst solid-state masers oscillators have been demonstrated at room temperature, continuous-wave amplification in these systems has only ever been realized at cryogenic temperatures. Here we report on a continuous-wave solid-state maser amplifier operating at room temperature. We achieve this feat using a practical setup that includes an ensemble of nitrogen-vacancy center spins in a diamond crystal, a strong permanent magnet and simple laser diode. We describe important amplifier characteristics including gain, bandwidth, compression power and noise temperature and discuss the prospects of realizing a room-temperature near-quantum-noise-limited amplifier with this system. Finally, we show that in a different mode of operation the spins can be used to reduce the microwave noise in an external circuit to cryogenic levels, all without the requirement for physical cooling.
@@ I Introduction
Maser amplification of short ($\sim 30~{}\mu$s) pulsed microwave signals was recently observed at room temperature using a spin system in an organic host [13], whilst a continuously operating diamond maser amplifier was realized at liquid nitrogen temperatures ($\sim 77$ K) [12]. However, the demonstration of a widely usable continuous-wave maser amplifier operating under ambient conditions remains a highly desirable objective.
@@
In this work we describe a continuous-wave maser amplifier based on a solid-state system that operates at room temperature. We use an ensemble of nitrogen-vacancy (NV) center spins in a bulk diamond crystal as the gain medium and couple this to a high quality factor microwave dielectric resonator. We achieve gains as high as 30 dB and a gain-bandwidth-product of up to 4.5 MHz. Our analysis shows that the intrinsic noise temperature of the spin gain medium is close to the quantum limit, whilst the dominant source of noise comes from resonator losses. We further demonstrate that the same device can be used to remove microwave noise in a circuit [16, 17, 18, 19, 20] by acting like a matched load with an effective microwave temperature of 66 K, without the need for any cryogenic cooling. We develop a modern quantum optics theoretical model [21] to describe the maser amplifier and microwave cooler system and obtain quantitative agreement with our measurement results. Finally, we discuss relatively straight-forward improvements that can be made to our system to push its noise performance towards the quantum limit. These results demonstrate that NV spin ensembles in diamond coupled to microwave resonators constitute a system with exceptional promise for performing ultra-low noise detection of microwave signals at room temperature.
@@ II.1 Maser amplifier design
In this work we utilize a high concentration NV- diamond bulk crystal as the gain medium. The energy level structure of the NV- triplet ground state is shown in Fig. 1a. When a static magnetic field $B_{\rm 0}$ is applied parallel to one of the four equivalent NV $\langle 111\rangle$ directions (see Fig. 1b), the $m_{\rm s}=\pm 1$ states of the NV- spins Zeeman split with a gradient of ±$\pm\gamma_{\rm e}$ (with $\gamma_{\rm e}/2\pi=28$ GHz/T), offset by a zero field splitting of $D/2\pi=2.87$ GHz [26, 27]. The NV- spins can be readily polarized into the ground triplet $m_{\rm s}=0$ state by optical pumping using green ($\sim$ 520 nm) light. At a sufficiently large $B_{\rm 0}$, the $m_{\rm s}=-1$ spin state energy falls below that of the the $m_{\rm s}=0$ state and the optical pumping results in a population inversion [9] (see Fig. 1c, lower), fulfilling a key condition for implementing a maser amplifier. For smaller $B_{\rm 0}$ fields where the $m_{\rm s}=0$ state is lowest in energy (see Fig. 1c, upper), the NV- spins absorb microwaves and can thus be used to reduce the noise present in a circuit [16, 17, 18, 19, 20].
@@ II.3.1 Gain
We proceed to operate the system as a maser amplifier by pumping the NV- spins with the laser diode at a power of $P_{\rm L}=1.61$ W, whilst simultaneously probing the resonator using a vector network analyzer (VNA), allowing us to measure the reflection coefficient $r^{+}$ (Eq. 2). We adjust the position of the permanent magnet poles to provide a field close to $B_{\rm 0}\approx 450$ mT, which brings the $m_{\rm s}=-1$ to $m_{s}=0$ spin transition near resonance with the dielectric resonator. We perform a fine sweep of the field using the Helmholtz coil and plot the power gain $G=20\log{|r^{+}|}$ as a function of $B_{\rm 0}$ in Fig. 3a. Away from resonance with the spins, we observe observe a dip in $G$ at $\omega_{\rm r}$, which is due to the internal resonator losses. As the spins come into resonance, they add energy to the system via stimulated emission, which compensates the internal losses in the resonator and transforms the dip in $G$ to a peak, signalling gain. We find a single gain peak, rather than the three that would be expected due to the hyperfine coupling of the NV- electronic spin with the nitrogen nuclear spin ($I=1$) [27]. This is because of the inhomogeneity of the $B_{\rm 0}$ field supplied by the permanent magnet at the given pole separation, which broadens the spin transition and produces a Gaussian-like profile (see Supplementary Information, Section LABEL:sup:subsec:spectra).
@@ II.4 Microwave cooling
Next we adjust the permanent magnet pole positions to set $B_{\rm 0}\approx 250$ mT, bringing the $m_{\rm s}=0$ to $m_{\rm s}=+1$ spin transition in resonance with the dielectric resonator. The NV- spins, which are now polarized in the ground state of this two-level subsystem, absorb microwave photons from the resonator (including thermal photons), which are then removed from the system via optical pumping. The removal of thermal noise photons lowers the effective microwave temperature of the mode [16, 17, 18, 19, 20], which can be transferred to an external microwave circuit.
@@
To find the cooling spin transition we perform a fine sweep of $B_{\rm 0}$ whilst probing the resonator in reflection, with the result shown in Fig. 5a. We observe avoided crossings in the measured spectrum, which is indicative of strong coupling between the resonator and hyperfine transitions of the NV- spins [43, 19]. We note that at this magnet pole separation we are able to partially resolve the hyperfine structure of the NV- system, indicating a higher field homogeneity than observed at the maser transition. Taking a line cut (dashed line in Fig. 5a) through the middle of this feature, we observe four dips in the magnitude response (see Fig. 5b, dark blue line), which are reproduced well with our model (see Eq. 2). A fit to the magnitude ($|r^{-}|$) and phase ($\angle r^{-}$) reflection data allows us to extract the ensemble coupling strength $g_{\rm ens}$, which is plot as an inset in Fig. 5b as a function of $P_{\rm L}$. We find excellent agreement with the $g_{\rm ens}$ extracted via the gain measurements in Section II.3.1.